OpenStudy (anonymous):

find the derivative of 14x^7 when x=1 to sqrt(x) a. 7x^3 b. 14x^(7/2) c. 7/4x^5-7/4 d. 28/3x^5

OpenStudy (anonymous):

integrate it or differentiate it?

OpenStudy (anonymous):

integrate it please

OpenStudy (anonymous):

integrate (14x^7) =(14/8)x^8 =(7/4)x^8 put x=1 (14/8)x^8 =(7/4) put sqrt(x) (14/8)x^8 =(14/8)x^4 =(7/4)x^4 ans = (7/4)x^4 -(7/4) =(14/8)(x^4) - (14/8)

OpenStudy (anonymous):

wow, thank you!

OpenStudy (anonymous):

you are welcome :D